# MECD: Unlocking Multi-Event Causal Discovery in Video Reasoning

### v2

@@ 4.1 VGCM: Video Granger Causality Model
$\begin{split}{\boldsymbol{F}}^{p}=\text{Enc}_{V}(\Phi_{pre}({\boldsymbol{V}}^{p})),~{}{\boldsymbol{O}}^{p}=\text{Dec}(\texttt{{Cat}}({\boldsymbol{F}}^{p},\text{Enc}_{C}({\boldsymbol{C}}^{p}))\\ {\boldsymbol{F}}^{m}_{k}=\text{Enc}_{V}(\Phi_{pre}({\boldsymbol{V}}^{m}_{k})),~{}{\boldsymbol{O}}^{m}_{k}=\text{Dec}(\texttt{{Cat}}({\boldsymbol{F}}^{m}_{k},\text{Enc}_{C}({\boldsymbol{C}}^{m}_{k}))\end{split}$ (1)
where $\text{Enc}_{V}$ and $\text{Enc}_{T}$ represent the encoder module of video clips and captions, respectively. Dec is a multi-modal decoder that shares weights for both paths. Cat indicates the concatenate operation. ${\boldsymbol{F}}^{p}$ and ${\boldsymbol{F}}^{m}_{k}$ are features after encoding, and ${\boldsymbol{O}}^{p}$ and ${\boldsymbol{O}}^{m}_{k}$ are the output features, which are then used for comparison of difference. Incorporating both visual and linguistic representations, the decoder conducts cross-modal reasoning and leverages the context from the unmasked premise events to posit a meaningful representation of the most likely explanatory result event.
@@ 5.1 Main results
As illustrated in Table LABEL:speed, we have assessed the inference speed of various models, with our VGCM achieving a swift 0.76 seconds per sample. The proposed method incurs an overhead of only 8.57% over the Videobert baseline. It is noteworthy that our inference speed is 3 to 6 times faster than that of all Video LLMs. The inference speed experiments were conducted on 1 NVIDIA A6000 GPU.
@@ 7 Acknowledgement
The paper is supported in part by the National Natural Science Foundation of China (No. 62325109, U21B2013).
@@ C.1 Modalities analysis of causality discovering
The MECD task leverages both video input and corresponding captions to uncover causality. In this experiment, our goal is to assess the degree of reliance on these two modalities in causal discovery. Typically, our MECD task captions consist of an average of 13.5 words per event, while each event comprises 50 frames. To investigate the influence of the text modality, we employ a masking strategy for the input information of the premise event, gradually increasing the masking ratio from 10% to 80%. The results are presented in Tab. 9.
The experimental results suggest a slight inclination towards the visual modality in the causality discovery task in Tab. 9, with a slightly lesser emphasis on text information. However, even with 80% masking of either modality, the results consistently outperform our strong baseline model, VGCM, underscoring the robust causal discovery capability of VGCM. Furthermore, we conducted experiments involving simultaneous masking of both modalities of information. Interestingly, we observe a noticeable decrease in accuracy compared to when only one modality is masked. This observation highlights the importance of jointly considering both modalities in the causality discovery task.

### v3

@@ 4.1 VGCM: Video Granger Causality Model
$\begin{split}{\boldsymbol{F}}^{p}=\text{Enc}_{V}(\Phi_{pre}({\boldsymbol{V}}^{p})),~{}{\boldsymbol{O}}^{p}=[\text{Dec}(\texttt{{Cat}}({\boldsymbol{F}}^{p},\text{Enc}_{C}({\boldsymbol{C}}^{p}))]_{N-1}\\ {\boldsymbol{F}}^{m}_{k}=\text{Enc}_{V}(\Phi_{pre}({\boldsymbol{V}}^{m}_{k})),~{}{\boldsymbol{O}}^{m}_{k}=[\text{Dec}(\texttt{{Cat}}({\boldsymbol{F}}^{m}_{k},\text{Enc}_{C}({\boldsymbol{C}}^{m}_{k}))]_{N-1}\end{split}$ (1)
where $\text{Enc}_{V}$ and $\text{Enc}_{T}$ represent the encoder module of video clips and captions, respectively. Dec is a multi-modal decoder that shares weights for both paths. Cat indicates the concatenate operation, and $[-]_{N-1}$ indicates the $(N-1)$-th slice at dimension 0. ${\boldsymbol{F}}^{p}$ and ${\boldsymbol{F}}^{m}_{k}$ are features after encoding, and ${\boldsymbol{O}}^{p}$ and ${\boldsymbol{O}}^{m}_{k}$ are the output features, which are then used for comparison of difference. Incorporating both visual and linguistic representations, the decoder conducts cross-modal reasoning and leverages the context from the unmasked premise events to posit a meaningful representation of the most likely explanatory result event.
@@ 5.1 Main results
As illustrated in Tab. LABEL:speed, we have assessed the inference speed of various models, with our VGCM achieving a swift 0.76 seconds per sample. The proposed method incurs an overhead of only 8.57% over the Videobert baseline. It is noteworthy that our inference speed is 3 to 6 times faster than that of all Video LLMs. The inference speed experiments were conducted on 1 NVIDIA A6000 GPU.
@@ 7 Acknowledgement
The paper is supported in part by the National Natural Science Foundation of China (No. 62325109, U21B2013) and the Lenovo Academic Collaboration Project.
@@ C.1 Modalities analysis of causality discovering
The MECD task employs both video input and corresponding captions to uncover causality. Our objective in this experiment is to evaluate the degree of reliance on these two modalities in causal discovery. Typically, each event in our MECD task consists of a textual input with an average of 13.5 words caption and a visual input of 50 frames.
To investigate the influence of the text modality, we employ a masking strategy for the input caption of the premise event, gradually increasing the masking ratio from 10% to 80%. The results presented in Tab. 9 indicate that our VGCM does not rely on the textual modality input; VGCM can also conduct causal discovery for videos without any captions.
In contrast, the experimental results suggest a more obvious performance decrease towards less visual modality input in the causality discovery task, as shown in Tab. 9. However, even with 80% masking of either modality, the results consistently outperform our strong baseline model, VideoLLaVA, underscoring the robust causal discovery capability of VGCM.
Furthermore, we conducted experiments involving simultaneous masking of both modalities of information. Interestingly, we observe a noticeable decrease in accuracy compared to when only one modality is masked. This observation highlights the importance of jointly considering both modalities in the causality discovery task.